OpenStudy (vshiroky):

Solve. Collect the variable terms on the left side of the inequality. 8x-2 is less than or equal to 3x+5

OpenStudy (vshiroky):

@johnweldon1993 I think it's 5x is less than or equal to 7

OpenStudy (anonymous):

x is less than or equal to \[\frac{ 7 }{ 5 }\]

OpenStudy (vshiroky):

is that because I have to just get the variable on the side?

OpenStudy (johnweldon1993):

You are correct @vshiroky But remember you need to solve for 'x' not 5x ... so you would divide both sides by 5 as crabby_patty has done...

OpenStudy (vshiroky):

ahh gotcha

OpenStudy (anonymous):

so u are correct basically....just don't forget the last step.

OpenStudy (johnweldon1993):

So you had \[\large 5x \le 7\] Just divide both sides by 5 to get \[\large x \le \frac{7}{5}\]

OpenStudy (vshiroky):

Makes sense now :) I'm posting another